# AI-driven Wireless Positioning: Fundamentals, Standards, State-of-the-art, and Challenges

### v1

@@ VI-B2 Geometric-knowledge-assisted Positioning
Another approach is to combine AI algorithms with geometric knowledge and spatial constraints to improve positioning accuracy. This includes tasks such as NLOS detection, TOA/angle estimation, and the integration of physical map constraints to enhance the understanding of spatial relationships between transmitters and receivers. When combined with AI algorithms, geometric knowledge allows these systems to adjust for variations in wireless environments or spatial shifts. For instance, in OTDOA-based positioning [333], the authors enhance accuracy by incorporating NLOS indicator with RSTD measurements as inputs to a neural network, which effectively mitigates NLOS problem and improves the positioning accuracy. Further extending this, the authors in [334] use hybrid delay and angular measurements in a neural network-based weighted least squares (WLS) framework to enhance performance. Another study [335] employs a deep variational learning method to estimate position-related parameters such as distance, TDOA, and AOA, which are subsequently used to calculate the position. In [336], the authors propose MLLoc, a ML-based positioning system that fuses GNSS and mobile network signals, using ML for stable TOA estimation to calculate positions accurately. In [337], the authors propose a classification-to-regression ANN model, leveraging class probabilities from classification to compute final positioning coordinates.
@@
In addition, positioning can be enhanced by utilizing environmental maps [338, 339], the correlation of user motion trajectories [340, 341], and the relationships between multiple BSs [342]. Incorporating prior physical map information, the authors in [338] propose a zero-shot learning framework for indoor localization using floor-plan images. A graph neural network is employed to model the relationships between APs and mobile devices for coarse localization, while floor-plan constraints refine positioning accuracy. The authors in [339] focus on multipath detection to reconstruct indoor environments. The study uses an ML model to predict dominant multipath components, identify virtual anchors, and build a generative channel model, significantly improving positioning accuracy under NLOS conditions. To address issues with outlier positioning estimates, the authors in [340] propose a context-aware localization technique that leverages historical trajectory information to refine the accuracy of anomalous points. In [341], the authors propose a DRL-based unsupervised wireless localization method, modeling localization as a Markov decision process and designing a reward-setting mechanism based on high RSS near APs for robust localization without retraining. In [342], the paper proposes a graph convolutional network based algorithm to model spatial relationships among multiple APs, extracting features from RSSI-based fingerprints for classification using an MLP.
@@ VI-C Channel Charting based Positioning
As mentioned above, wireless-knowledge-assisted positioning builds upon fingerprint-based positioning by leveraging wireless knowledge to further enhance localization performance and reduce the reliance on extensive data collection. Building on this foundation, Channel charting based positioning eliminates the need for external RPsor exhaustive datasets by directly learning latent spatial relationships from CSI data, enabling relative or pseudo-positioning. This approach allows channel charting to be highly adaptable to dynamic environments while significantly reducing operational overhead, making it a promising alternative to traditional localization methods in rapidly changing wireless scenarios.
@@
Numerous studies have investigated the potential of channel charting to enhance wireless positioning performance [343, 344, 345, 346, 166, 347, 348, 349, 160, 350]. The authors in [343] first introduce channel charting for positioning by proposing a unified Siamese network architecture for CSI-based localization. Their framework supports supervised, semi-supervised, and unsupervised scenarios, leveraging Sammon’s mapping extension and side information to achieve accurate positioning in both LOS and NLOS channels with minimal CSI measurements. Building on this, graph-based approaches are further applied to channel charting and wireless positioning. For example, the authors in [344] propose a semi-supervised graph-based channel charting framework for 5G localization that utilizes distributed CSI, side-information, and constrained manifold learning to construct a 2D channel chart, achieving 5.6 m localization accuracy with minimal labeled samples. In [345], a comparative study between classical model-based localization methods and channel charting highlight the limitations of channel charting in achieving global geometric accuracy. To address this, augmented channel charting is proposed by integrating model-based localization information into channel charting training, thereby improving overall performance and surpassing classical methods on evaluated datasets. Leveraging temporal correlations during sampling, the authors in [346] introduce a reference-free channel charting framework that incorporates velocity information and topological maps to transform relative charts into real-world coordinates. Further enhancements are made in [166], where a novel dissimilarity metric is introduced, incorporating angular-domain information and a deep learning-based metric. Additionally, metric fusion is proposed to integrate temporal and CSI similarities, demonstrating superior performance in sub-6 GHz massive MIMO scenarios, even under NLOS conditions. In [347], a Doppler-based loss function for channel charting is introduced, requiring only frequency synchronization to enable localization with minimal assumptions. The authors in [349] develop a geodesic distance-based metric for channel charting using synchronized CSI measurements, utilizing a Siamese network to learn global geometry for localization, which outperform traditional fingerprinting methods in real-world 5G and UWB systems. Additionally, in [160], the paper presents a multi-point channel charting approach for multi-gateway LoRa networks, applying t-SNE and k-means clustering on received power vectors to map spatial geometry and improve IoT device localization. Lastly, privacy concerns in channel charting are systematically examined in [350], focusing on user and vendor privacy risks associated with pseudo-locations and raw CSI exposure.
@@ VII Datasets for Wireless Positioning
In recent years, several datasets have been proposed in the field of wireless positioning, including the xG-Loc dataset [351], DeepMIMO dataset [352], Wireless AI Research Dataset (WAIR-D) [353], DataAI-6G dataset [354] and the ViWi dataset [355]. However, the above datasets do not conform to the clustered delay line (CDL) channel model [356] recommended by 3GPP. Since CDL channel model is more suitable for link-level and system-level simulation and for the comprehensiveness of the study, we also examine datasets based on CDL channel model.
VII-1 xG-Loc
@@
The xG-Loc dataset [351] is the first open dataset explicitly designed for localization algorithms and services, fully compliant with 3GPP technical reports and specifications. This dataset provides a standardized and comprehensive resource for evaluating and benchmarking localization solutions across diverse scenarios. xG-Loc is structured into 28 compressed directories, representing 28 unique configurations of 3GPP-standardized scenarios, bandwidths, and central frequencies. These configurations span multiple frequency ranges, including FR1 (microwaves), FR2 (millimeter waves), and FR3 (upper mid-band), thereby enabling comprehensive analysis across various operational conditions. The scenarios considered in the xG-Loc dataset align with 3GPP 38.901 [356].
Each directory contains a rich set of data, including PRS and SRS, measurement data, analytics, position estimates, and channel quality indicators (CQIs). The dataset is organized in text and JSON files to facilitate flexible usage and compatibility with various ML frameworks and localization algorithms.
VII-2 5G-NR-data-generator
The 5G-NR-data-generator [357] is a generalized channel dataset generator that adopts the CDL channel model of the 5G NR standard [356]. This generator allows users to customize various channel parameters to meet specific needs and supports the generation of MIMO channels.
Figure 10: Dataset generation process of 5G-NR-data-generator.
The data generation process is shown in Fig. 10. The configurable channel parameters $S^{\prime}$ include a variety of settings, such as the number of RBs, subcarriers, bandwidth, the number of BS and UE antennas, and the channel center frequency. Based on the user-defined parameters $S^{\prime}$ or $S_{u}^{\prime}$, the 5G-NR-data-generator performs ray-tracing simulations to model the propagation environment. The ray-tracing simulation determines the propagation paths from the BS to the UE by utilizing the geometric surface data provided in the map files.
These map files and terrain data are sourced from the open-source Open Street Map (OSM) project [358]. Using the terrain and map data, the simulation calculates key channel characteristics, including propagation delays, AOA, AOD, LOS or NLOS labels, and path gains. These outputs are then combined to construct the final channel model.
VII-3 DeepMIMO
@@
DeepMIMO [352], developed by researchers from Arizona State University, is a versatile and generic deep learning dataset designed specifically for mmWave and massive MIMO applications. The dataset provides a comprehensive framework for generating channel datasets tailored to various tasks in wireless communication.
The channels in the DeepMIMO dataset are constructed using accurate ray-tracing data obtained from the Wireless InSite simulator [359]. This ray-tracing simulation captures the dependence of the wireless channels on the geometry and materials of the environment, as well as the spatial locations of the transmitter and receiver. This feature is crucial for ML applications in mmWave and massive MIMO systems, as it ensures the dataset realistically reflects the physical characteristics of the propagation environment.
Based on DeepMIMO, researchers can customize the dataset by configuring a set of parameters to suit their specific ML tasks. These parameters control various system and channel elements, including the number of antennas, subcarriers, channel paths, and other factors. Therefore, the DeepMIMO dataset is defined by two primary components: a scenario and a set of customizable parameters. The process of using the DeepMIMO dataset generation framework is straightforward, as illustrated in Fig. 11.
VII-4 WAIR-D
The WAIR-D [353] is developed by researchers from Huawei and Zhejiang University to provide a wireless dataset that replicates various environments closely resembling real-world conditions. The dataset is generated using a 3D ray-tracing simulator, PyLayers [360], in conjunction with the OSM [358]. PyLayers utilizes environmental information, such as maps containing the locations of BSs and UEs, to calculate radio propagation paths for each radio link between BSs and UEs. The use of OSM provides details about real-world building layouts and street directions, enhancing the authenticity and diversity of the generated data.
WAIR-D offers flexibility in generating channels with customizable wireless system parameters, supporting five carrier frequencies, i.e., 2.6 GHz, 6 GHz, 28 GHz, 60 GHz, and 100 GHz, enabling applications in both sub-6 GHz and mmWave scenarios. Users can configure system bandwidth, numerology, and the number of antennas, tailoring the dataset to specific requirements. WAIR-D includes two scenarios: one with 10,000 environments featuring sparsely distributed UEs for wide coverage and another with 100 environments focusing on densely deployed UEs for high-density urban settings.
The dataset provides a rich array of details about the propagation paths and the environment, including path delays, AOA, AOD, and other channel characteristics. The data is stored in specific file formats to facilitate its use in various wireless research applications.
VII-5 DataAI-6G
@@
The DataAI-6G dataset [354], developed by researchers from Beijing University of Posts and Telecommunications and the China Mobile Research Institution, is specifically designed for AI-6G research. A key feature of this dataset, and its major advantage over the DeepMIMO and WAIR-D datasets, is its incorporation of spatial non-stationary features and high-mobility characteristics. With the adoption of ultra-massive MIMO and the increasing demand for high-mobility communications, these features are becoming more prominent, making it essential to include them in the dataset. This inclusion significantly enhances positioning accuracy in corresponding scenarios.
A notable advantage of the DataAI-6G dataset [354] is its generic framework, which allows researchers to configure parameters according to their specific requirements. This framework enables users to input raw channel parameters and generate customized datasets tailored to their research needs. The dataset provides detailed channel information obtained from the Wireless InSite ray-tracing simulator [359], including parameters such as AOD, AOA, delay, phase, power of each path, path loss between each pair of transceiver antennas, and features such as spatial non-stationarity and mobility.
Therefore, the dataset is particularly well-suited for validating mobility in cross-band scenarios, making it highly applicable to advanced 6G research. Furthermore, it is fully compatible with ML and AI algorithms, enabling researchers to explore innovative solutions for mobility and non-stationary environments in 6G systems.
VII-6 ViWi
The ViWi dataset [355] is designed specifically for vision-aided wireless communications research. It serves as a parametric, systematic, and scalable data generation framework, leveraging advanced 3D modeling and ray-tracing software to produce high-fidelity synthetic wireless and vision data samples for identical scenes. By combining vision and wireless data, the dataset facilitates research at the intersection of these domains, supporting innovative approaches to wireless communication and positioning.
The ViWi dataset includes diverse signal characteristics such as AOD, AOA, path gains, and images. It also provides detailed user location data, enabling researchers to explore localization and communication techniques in visually rich environments. The dataset demonstrates flexibility and diversity through its parametric customization capabilities. Visual raw data is stored in JPEG and MAT formats, while wireless raw data comprises three MAT files per transmitter, providing a comprehensive and well-structured resource for research. The dataset generation process consists of three stages, as illustrated in Fig. 12:

### v2

@@ VI-B2 Geometric-knowledge-assisted Positioning
TABLE XIV: Categorization of Geometric-Knowledge-Assisted Positioning Methods
Category Methods & References
Knowledge from geometric information Based on distance and angle measurements [330, 331, 332, 333], LOS/NLOS classification [331, 150], and motion trajectory analysis [334, 335]
Knowledge from environmental priors Based on physical maps [336, 337, 338] and BS deployment [339]
Another effective approach to improving positioning accuracy is to combine AI algorithms with geometric knowledge and spatial constraints. When integrated with AI models, geometric knowledge enables the system to better adapt to variations in wireless environments and spatial dynamics. Specifically, this category leverages either geometric information or environmental priors to assist positioning. Table. XIV presents the classification and representative studies of geometric-knowledge-assisted positioning methods.
@@
From the geometric perspective, methods utilize distance and angle measurements [330, 331, 332, 333], LOS/NLOS classification [331, 150], and motion trajectory analysis [334, 335] to infer accurate locations. For instance, in OTDOA-based positioning [150], the authors enhance accuracy by incorporating NLOS indicator with RSTD measurements as inputs to a neural network, which effectively mitigates the NLOS problem and improves the positioning accuracy. Further extending this, the authors in [331] use hybrid delay and angular measurements in a neural network-based weighted least squares (WLS) framework to enhance performance. Another study [330] employs a deep variational learning method to estimate position-related parameters such as distance, TDOA, and AOA, which are subsequently used to calculate the position. To address issues with outlier positioning estimates, the authors in [334] propose a context-aware localization technique that leverages historical trajectory information to refine the accuracy of anomalous points. In [335], the authors propose a DRL-based unsupervised wireless localization method, modeling localization as a Markov decision process and designing a reward-setting mechanism based on high RSS near APs for robust localization without retraining.
@@
From the perspective of environmental and configuration priors, techniques leverage physical maps [336, 337, 338] (such as floor plans) and BS deployment information [339] to enhance localization. For example, [336] introduces a zero-shot learning framework for indoor positioning using floor plan images. A graph neural network is employed to model the relationships between APs and devices for coarse localization, while floor-plan constraints refine positioning accuracy. The authors in [337] focus on multipath detection to reconstruct indoor environments. The study uses an ML model to predict dominant multipath components, identify virtual anchors, and build a generative channel model, significantly improving positioning accuracy under NLOS conditions. In [339], the paper proposes a graph convolutional network to model spatial relationships among multiple APs, extracting features from RSSI-based fingerprints.
@@ VI-C Channel Charting based Positioning
As mentioned above, wireless-knowledge-assisted positioning builds upon fingerprint-based positioning by leveraging wireless knowledge to further enhance localization performance and reduce the reliance on extensive data collection. Building on this foundation, Channel charting-based positioning eliminates the need for external RPs or exhaustive datasets by directly learning latent spatial relationships from CSI data, enabling relative or pseudo-positioning. This approach allows channel charting to be highly adaptable to dynamic environments while significantly reducing operational overhead, making it a promising alternative to traditional localization methods in rapidly changing wireless scenarios.
@@
Numerous studies have investigated the potential of channel charting to enhance wireless positioning performance [340, 341, 342, 343, 129, 344, 345, 346, 123, 347]. The authors in [340] first introduce channel charting for positioning by proposing a unified Siamese network architecture for CSI-based localization. Their framework supports supervised, semi-supervised, and unsupervised scenarios, leveraging Sammon’s mapping extension and side information to achieve accurate positioning in both LOS and NLOS channels with minimal CSI measurements. Building on this, graph-based approaches are further applied to channel charting and wireless positioning. For example, the authors in [341] propose a semi-supervised graph-based channel charting framework for 5G localization that utilizes distributed CSI, side-information, and constrained manifold learning to construct a 2D channel chart, achieving 5.6 m localization accuracy with minimal labeled samples. In [342], a comparative study between classical model-based localization methods and channel charting highlight the limitations of channel charting in achieving global geometric accuracy. To address this, augmented channel charting is proposed by integrating model-based localization information into channel charting training, thereby improving overall performance and surpassing classical methods on evaluated datasets. Leveraging temporal correlations during sampling, the authors in [343] introduce a reference-free channel charting framework that incorporates velocity information and topological maps to transform relative charts into real-world coordinates. Further enhancements are made in [129], where a novel dissimilarity metric is introduced, incorporating angular-domain information and a deep learning-based metric. Additionally, metric fusion is proposed to integrate temporal and CSI similarities, demonstrating superior performance in sub-6 GHz massive MIMO scenarios, even under NLOS conditions. In [344], a Doppler-based loss function for channel charting is introduced, requiring only frequency synchronization to enable localization with minimal assumptions. The authors in [346] develop a geodesic distance-based metric for channel charting using synchronized CSI measurements, utilizing a Siamese network to learn global geometry for localization, which outperforms traditional fingerprinting methods in real-world 5G and UWB systems. Additionally, in [123], the paper presents a multi-point channel charting approach for multi-gateway LoRa networks, applying t-SNE and k-means clustering on received power vectors to map spatial geometry and improve IoT device localization. Lastly, privacy concerns in channel charting are systematically examined in [347], focusing on user and vendor privacy risks associated with pseudo-locations and raw CSI exposure.
VI-D Lessons Learned
From our comprehensive analysis of direct AI/ML positioning techniques, several important observations emerge. Fingerprint-based positioning, while capable of delivering high accuracy through advanced deep learning architectures such as CNNs, LSTMs, and Transformers, is heavily reliant on dense labeled data and frequent updates, making it costly and less scalable in dynamic or large-scale scenarios. In contrast, knowledge-assisted AI positioning addresses data inefficiency and generalization issues by embedding domain knowledge (wireless channel characteristics and geometric constraints) into the learning process, thereby improving robustness and reducing dependence on exhaustive measurements. Channel charting offers a compelling alternative by learning latent spatial structures directly from CSI data without the need for labeled positions, making it highly scalable and label-efficient. However, challenges remain in achieving global geometric accuracy and ensuring interpretability. Therefore, future research must focus on how to effectively exploit both wireless and geometric knowledge in combination with cross-domain adaptation techniques, in order to enable accurate, robust, and environment-adaptive positioning. Lastly, the increasing emphasis on lightweight and interpretable models reflects the growing demand for real-time, resource-efficient localization solutions in practical deployments. These findings highlight that future advances in AI-driven positioning will likely hinge on the integration of domain knowledge, data-efficient learning, and scalable system design.
@@ VII Datasets for Wireless Positioning
TABLE XV: Comparison of Wireless Positioning Datasets
Dataset Name Channel Model Type Open Source Type Mobility Support Parameter Configuration Limitation
xG-Loc CDL channel model (3GPP 38.901) Simulated dataset No Adjustable bandwidth (5–400 MHz), frequency (FR1/FR2/FR3), and scenario type (UMi, UMa, RMa, InF-DH, and InF-SH) No measured data; static channels only.
MaMIMO Real measurement Measured dataset (252,004 samples) Yes 2.61 GHz frequency, 20 MHz bandwidth, four antenna array topologies (URA LOS/NLOS, ULA LOS, DIS LOS) Limited to indoor environments; no high-speed mobility support.
DeepMIMO Ray tracing Channel simulator No Antenna array structure, OFDM subcarriers, bandwidth, frequency, scenario selection, BS/UE range, number of paths Fixed sampling points; cannot simulate dynamic channels.
5G-NR-data-generator CDL model + ray tracing Channel simulator No RBs, subcarriers, bandwidth, BS/UE antenna count, carrier frequency, delay/angle spread Single fixed scene; limited environment diversity
WAIR-D Ray tracing Channel simulator No Carrier frequencies, antenna configuration, bandwidth, OFDM settings Fixed environments; no dynamic channel or mobility modeling.
DataAI-6G Ray tracing Channel simulator Yes Adjustable frequency band, bandwidth, antenna config, user mobility speed/direction, sampling interval Only one outdoor street scenario; lacks environment diversity
ViWi Ray tracing + 3D visual simulation Channel/vision simulator No BS/UE positions, antenna config, frequency, trajectory, camera setup, visual resolution Synthetic data only; limited scenarios; no Doppler modeling.
VII-A Dataset Descriptions
@@
In recent years, several datasets have been proposed in the field of wireless positioning, including the xG-Loc [348], MaMIMO [349], DeepMIMO, [350], 5G-NR-data-generator [351], Wireless AI Research Dataset (WAIR-D) [352], DataAI-6G [353], and the ViWi [354]. However, the above datasets do not conform to the clustered delay line (CDL) channel model [355] recommended by 3GPP. Since the CDL channel model is more suitable for link-level and system-level simulation and for the comprehensiveness of the study, we also examine datasets based on the CDL channel model. A comparison of wireless positioning datasets is shown in Table. XV.
VII-A1 xG-Loc
@@
The xG-Loc dataset [348] is an open dataset explicitly designed for localization algorithms and services, fully compliant with 3GPP technical reports and specifications. This dataset provides a standardized and comprehensive resource for evaluating and benchmarking localization solutions across diverse scenarios, conforming to the channel models defined in 3GPP TR 38.901 [355]. xG-Loc is structured into 28 compressed directories, representing 28 unique configurations of 3GPP-standardized scenarios, bandwidths, and central frequencies. These configurations span multiple frequency ranges, including FR1 (microwaves), FR2 (millimeter waves), and FR3 (upper mid-band), as well as a variety of deployment scenarios, such as Indoor Factory Dense-High (InF-DH), Indoor Factory Sparse-High (InF-SH), Indoor Open Office (IOO), Urban Microcell (UMi), Urban Macrocell (UMa), Rural Macro (RMa). In addition, the dataset provides PRS, SRS, and other channel-related measurements. It enables a wide range of tasks, including distance estimation, angle estimation, position estimation, and wireless channel quality evaluation.
VII-A2 MaMIMO
MaMIMO [349] is an open-access indoor CSI dataset, providing 252,004 high-precision CSI samples collected using the 64-antenna KU Leuven Massive MIMO testbed. Measurements were conducted across four antenna topologies: Uniform Rectangular Array (URA) in both LOS and NLOS conditions, Uniform Linear Array (ULA), and Distributed ULA (DIS), each within a controlled indoor office environment. CSI was recorded over a 20 MHz bandwidth at 2.61 GHz using 100 subcarriers, with user equipment moved across a 1.25 m × 1.25 m grid via CNC-controlled XY-tables. This setup ensured sub-millimeter accuracy in positional labeling. The resulting dataset enables benchmarking of positioning algorithms, multipath component analysis, and precoder visualization.
VII-A3 5G-NR-data-generator
The 5G-NR-data-generator [351] is a generalized channel dataset generator that adopts the CDL channel model defined in the 5G NR standard [355]. This generator allows users to customize various channel parameters to meet specific needs and supports the generation of MIMO channel matrices. The configurable channel parameters include a variety of settings, such as the number of RBs, subcarriers, bandwidth, number of BS/UE antennas, carrier frequency, delay spread, and angle spread. The generator performs ray-tracing-based simulations to model the propagation environment using external tools. However, the current version only supports a single predefined scenario and does not allow flexible environment modeling or dynamic user movement.
VII-A4 DeepMIMO
@@
DeepMIMO [350] provides a comprehensive framework for generating channel datasets tailored to various tasks in wireless communication. The channels in the DeepMIMO dataset are constructed using accurate ray-tracing data obtained from the Wireless InSite simulator [356]. This ray-tracing simulation captures the dependence of the wireless channels on the geometry and materials of the environment, as well as the spatial locations of the transmitter and receiver. This feature is crucial for ML applications in mmWave and massive-MIMO systems, as it ensures the dataset realistically reflects the physical characteristics of the propagation environment. DeepMIMO enables researchers to flexibly configure their simulation scenarios and wireless parameters prior to dataset generation. These parameters include antenna array structure, OFDM subcarrier configuration, bandwidth, frequency band, scenario selection, user/BS activation range, and the number of paths. Therefore, the DeepMIMO dataset is defined by two primary components: a scenario and a set of parameters. However, the channel generation in DeepMIMO is based on pre-collected ray-tracing paths at fixed sampling points, offering limited flexibility in sampling location configuration. Moreover, dynamic channels caused by user movement cannot be simulated due to the lack of channel temporal modeling. The process of using the DeepMIMO dataset generation framework is illustrated in Fig. 8.
VII-A5 WAIR-D
The WAIR-D [352] is developed by researchers from Huawei and Zhejiang University to provide a wireless dataset that replicates various environments closely resembling real-world conditions. The dataset is generated using a 3D ray-tracing simulator, PyLayers [357], in conjunction with the OSM [358]. WAIR-D includes two scenarios: 1). Scenario 1 consists of 10,000 environments, each with 5 base stations and 30 sparsely distributed user equipments (UEs); 2). Scenario 2 contains 100 environments, each with 1 base station and 10,000 densely dropped UEs. WAIR-D supports five carrier frequencies, i.e., 2.6 GHz, 6 GHz, 28 GHz, 60 GHz, and 100 GHz. Users can configure antenna configuration, bandwidth, and OFDM subcarrier settings to generate channel data. Similar to DeepMIMO, WAIR-D is based on static ray-tracing results and does not support flexible sampling or dynamic channel modeling caused by user movement.
VII-A6 DataAI-6G
@@
The DataAI-6G dataset [353], developed by researchers from Beijing University of Posts and Telecommunications and the China Mobile Research Institution, is specifically designed for AI-6G research. Compared to existing datasets such as DeepMIMO and WAIR-D, DataAI-6G introduces key advancements by incorporating spatially non-stationary channel features and modeling high-mobility scenarios, including Doppler effects and time-varying multipath propagation. The dataset allows users to configure a wide range of system parameters, such as carrier frequency (uplink and downlink), bandwidth, antenna array configuration, user mobility paths and speeds, number of users, and OFDM settings. However, its current version is limited to a single outdoor street scenario, lacking environmental diversity across indoor, suburban, or industrial settings.
VII-A7 ViWi
The ViWi dataset [354] is designed specifically for vision-aided wireless communications research. It serves as a parametric, systematic, and scalable data generation framework, leveraging advanced 3D modeling and ray-tracing software to produce high-fidelity synthetic wireless and vision data samples for identical scenes. By combining vision and wireless data, the dataset facilitates research at the intersection of these domains, supporting innovative approaches to wireless communication and positioning. The dataset includes wireless parameters such as AOD, path gains, and detailed user location, along with corresponding RGB images, enabling joint research in communication and localization. ViWi supports customizable scene configurations and user trajectories but is limited to synthetic data and does not model Doppler effects or continuous channel dynamics. The dataset generation process consists of three stages, as illustrated in Fig. 9: